# Pre-Training Representations of Binary Code Using Contrastive Learning

### v4

@@ 4 Experimental Results
Figure 6: Embedding analysis of pre-trained models for binary code. In each subfigure, the red and orange dots represent the embeddings of the keymap function and key record function binary codes, respectively. The lower part shows their source and binary code. RoBERTa, CodeBERT, and GraphCodeBERT generate separate embeddings for similar semantics, while ContraBin (PCL) and ContraBin cluster them together.
@@ 4.1 Embedding Analysis
TABLE III: Quantitative evaluation of ContraBin on binary functional algorithm classification of POJ104 using mean of average precision (MAP), recall (MAR) and F1 score (MAF1).
@@
RoBERTa [13] 38.29 28.57 32.72
CodeBERT [12] 39.99 (+4.44%) 29.94 (+4.80%) 34.24 (+4.64%)
GraphCodeBERT [9] 39.48 (+3.11%) 29.38 (+2.95%) 33.69 (+2.96%)
@@
TABLE IV: Quantitative evaluation of ContraBin and multiple pre-trained methods on binary function name recovery of DIRE using accuracy, mean of average prevision (MAP), mean of average recall (MAR), and mean of average F1 score (MAF) as four key criteria.
@@
RoBERTa [13] 29.41 28.41 25.00 26.59
CodeBERT [12] 24.94 (-15.20%) 23.25 (-18.15%) 20.82 (-16.70%) 21.97 (-17.39%)
GraphCodeBERT [9] 25.95 (-11.76%) 27.59 (-2.87%) 22.52 (-9.91%) 24.80 (-6.74%)
@@ 4.2 RQ1: Binary Code Analysis
After analyzing embedding of binary code using all pre-trained models, we start qualitative binary code analysis of ContraBin by two downstream tasks: algorithmec functionality classification and binary functiona name recovery.
@@
To evaluate the performance of ContraBin in this downstream task, we adopt mean of average prevision (MAP), mean of average recall (MAR), and mean of average F1 score (MAF) as three key criteria. We consider only C code that can be compiled to binaries — while most programs compiled successfully, we exclude those that could not be compiled, and report MAP, MAR, and MAF across all 104 classes. For this task and the three subsequent tasks, we used the results of RoBERTa as a baseline for comparison.
Table III presents the quantitative evaluation of ContraBin against several pre-trained methods. CodeBERT and GraphCodeBERT shows comparable performance to RoBERTa on all three metrics, with a performance improvement of less than 5%. However, ContraBin demonstrates a substantial improvement in MAP, MAR, and MAF1 by 14.34%, 18.55%, and 16.71%, respectively. We note that the performance of ContraBin (PCL) is lower than that of RoBERTa. This can be partially attributed to the use of contrastive learning over three different modalities, which can result in the binary code embedding becoming overfitted to the exact embedding of either source code or comments, making the model less effective.
@@
The overall performance, as shown in Table IV, revealed that fine-tuning CodeBERT and GraphCodeBERT resulted in a decrease in performance, making them less effective than the baseline RoBERTa model in all metrics by up to approximately 20%. This demonstrated that current models are unable to generalize when the domain of downstream applications for binary code analysis shifts and may even become ineffective. On the other hand, ContraBin still demonstrated considerable improvement over all three metrics by 13.6%, 8.39%, and 10.58%, respectively, demonstrating the generalizability of our pretrained model on different downstream tasks.
TABLE V: Quantitative evaluation of ContraBin and multiple pre-trained methods on binary code summarization of AnghaBench using BLEU-4, GLUE-4, ROUGE-2, and exact match (xMatch). Performance differences are calculated relative to GraphCodeBERT.
Approaches BLEU-4 GLUE-4 ROUGE-2 xMatch
RoBERTa [13] 24.30 (-25.05%) 25.55 (-22.45%) 30.16 (-23.52%) 4.95 (-37.41%)
CodeBERT [12] 32.07 (-1.11%) 32.80 (-0.43%) 38.83 (-1.57%) 7.91 (0.00%)
GraphCodeBERT [9] 32.43 (Base) 32.94 (Base) 39.45 (Base) 7.91 (Base)
@@
TABLE VI: Quantitative evaluation of ContraBin on binary reverse engineering of AnghaBench using BLEU-4, GLUE-4 and exact match (xMatch).
Approaches BLEU-4 GLUE-4 xMatch
RoBERTa [13] 69.72 68.29 23.52
CodeBERT [12] 70.37 (+0.94%) 67.27 (-1.50%) 23.41 (-0.47%)
GraphCodeBERT [9] 71.56 (+2.64%) 69.88 (+2.32%) 24.73 (+5.14%)
@@ 4.3 RQ2: Binary Code Comprehension
Result for AnghaBench (Summarization) For the binary code summarization task, we evaluate all models using four widely adopted metrics: BLEU-4 [27], GLUE-4 [28], ROUGE-2 [29], and Exact Match (xMatch). As shown in Table V, ContraBin achieves consistent improvements across all metrics compared to the baseline method, GraphCodeBERT. Specifically, ContraBin shows improvements of 5.95% in BLEU-4, 5.71% in GLUE-4, 4.43% in ROUGE-2, and 18.08% in xMatch over GraphCodeBERT, showcasing its superior capability in binary code summarization.
To further highlight the effectiveness of ContraBin, we present a case study in Figure 7. In the first example, ContraBin successfully captures the functionality of the binary program, while other models diverge to irrelevant objectives, such as ”cpu,” ”file,” or ”lock.” In the second example, other models struggle with summarizing longer binary code. For instance, ContraBin (PCL) generates incomplete sentences, whereas RoBERTa, CodeBERT, and GraphCodeBERT produce either incorrect or overly general summaries. These cases demonstrate ContraBin’s ability to maintain long-term semantic consistency and robustness in binary code summarization, enabled by its novel binary embeddings.
Result for AnghaBench (Reverse Engineering) For the binary code reverse engineering task, we use similar metrics, except ROUGE, which is designed for summarization only, to evaluate the performance of ContraBin on direct code translation between source code and binary code. As shown in Table VI, ContraBin improves the baseline method by 2.04% in BLEU, 3.10% in GLEU, and 7.01% in xMatch. Compared with GraphCodeBERT, which outperforms ContraBin in BLEU, ContraBin outperforms on the other two metrics (GLUE and xMatch).
@@ 5 Ablation Studies and Model Insights
TABLE VII: Ablation studies of ContraBin on binary functional algorithm classification of POJ104 using mean of average precision (MAP), recall (MAR), and F1 score (MAF1). Differences are noted in points (pt).
Ablations MAP MAR MAF1
Human-written comments 3.11 (-40.67 pt) 3.01 (-30.86 pt) 3.06 (-35.13 pt)
Comment 42.76 (-1.02 pt) 33.23 (-0.64 pt) 37.40 (-0.79 pt)
Anchor 40.96 (-2.82 pt) 31.07 (-2.80 pt) 35.34 (-2.85 pt)
Interpolations 37.03 (-6.75 pt) 26.99 (-6.88 pt) 31.22 (-6.97 pt)
None (ContraBin) 43.78 33.87 38.19
@@ 5.1 Ablation Studies
We assess each component of our model through ablation studies. Specifically, we remove three conditions to test the efficiency of each design element: (1) We remove comments as one of the modalities during training, (2) we replace the anchored encoder with a learnable one so that it can be updated during training, and (3) we eliminate both linear and non-linear interpolation from our model design. For the pretraining process, we adhere to the original hyperparameters, training the first and second ablations for 30 epochs and the third ablation for 10 epochs. For the third ablation, we ensure that the model is converged by the end of pretraining. We test the pretrained models on the binary functional algorithm classification of POJ140, using the same hyperparameters. The results are shown in Table VII.
@@ 5.2 Comparison of Human-Written and LLM-Generated Comments in Pretraining
Performance comparison The experiments reveal a striking decrease in performance when ContraBin is trained with human-written comments compared to LLM-generated comments. Specifically, Table VII shows that the use of human-written comments results in MAP dropping from 43.78 to 3.11, MAR from 33.87 to 3.01, and MAF1 from 38.19 to 3.06. These reductions of 40.67 pt, 30.86 pt, and 35.13 pt, respectively, highlight the stark contrast in effectiveness between the two types of comments.
@@ 5.5 Extensibility to Contemporary LLMs
The evolving landscape of LLMs necessitates the adaptation of existing methodologies to newer, more advanced architectures. As LLMs like T5 [30], CodeT5 [24] and its derivatives become increasingly prominent, it is crucial to ensure that techniques initially developed for models like RoBERTa can be effectively extended to these contemporary frameworks. This section outlines the key adaptations required to transition from RoBERTa’s encoder-only architecture to T5’s encoder-decoder structure.
@@
TABLE VIII: Performance Comparison of ContraBin and CodeT5 with T5 on Binary Functional Algorithm Classification for POJ104 (MAP, MAR, MAF1).
@@
As shown in Table VIII, the ContraBin (CodeT5) model with our pretraining method achieves significant improvements over CodeT5. Specifically, our model achieves a MAP of 30.1, which is a 4.45% improvement over CodeT5, a MAR of 22.16 (+2.07%), and a MAF1 of 25.53 (+3.11%). These results underscore the effectiveness of our training methodology in enhancing the model’s ability to comprehend and process binary code, demonstrating the adaptability and strength of the T5 architecture.
@@ 6 Related Work
Binary code analysis Binary code analysis plays an important role in the bigger domain of software analysis and maintenance. For example, tracing execution can help the analysis of the functionality algorithms from binaries [3], reusing profile information can speed up the similarity comparison of frequently executed core code [2], incorporating sequences of system calls can help detect software vulnerabilities [5], and combining machine and binary-interface descriptions can assist software reverse engineering [4]. Traditionally, in programming language and software engineering, researchers have presented various work based on static and dynamic program analysis techniques. For instance, BYTEWEIGHT [6] automatically classifies algorithms by functionality from binaries, BINSEC [31, 7] formalizes low-level regions of code to extract semantics on the programs, and BinSide [8] uses intermediate representations to conduct cross-platform static analysis.
@@
Neural code models In recent years, neural code models have attracted the attention of researchers in software engineering and security. With the assistance of AI [32], it can improve many source code analysis tasks, such as neural clone detection, neural code similarity comparison, neural code search, etc [33, 34, 35, 36, 37]. For example, SySeVR combines syntax, semantics, and neural vector representation to detect software vulnerabilities [38], InnerEye adopts techniques in Natural Language Processing (NLP) to compare between samples of binary code [39], GNN-BPE [40] applies Graph Neural Networks (GNN) on binary code by fusing the semantics of Control Flow Graphs (CFG), Data Flow Graphs (DFG), and call graphs, and Bin2Vec [41] learns binary code representation via Graph Convolutional Networks (GCN). Compiler chain detection has also emerged as a promising application of machine learning in binary analysis, enabling the identification of the toolchains used to generate binary executables and aiding tasks like provenance analysis and malware detection [42, 43]. Additionally, the survey by Marcelli et al. [44] provides a detailed review of how machine learning techniques are solving the binary function similarity problem, which complements the above-discussed approaches. However, these methods involve complex aggregation of additional binary code representations and have limited generalizability across tasks.
@@
Large-scale pre-trained embeddings With the recent success of large-scale pre-trained embedding in AI, research has been conducted on applying similar approaches to code analysis [33, 45, 46, 47]. In particular, RoBERTa [13] serves as the baseline, CodeBERT [12] incorporate code generation and identification into training process, GraphCodeBERT [9] combines Abstract Syntax Tree (AST) to further improve embedding quality. There are also many recent models designed for source code understanding, including CodeTrans [48] and CoTexT [49] that incorporate different language modalities and multi sub-tasks for better code representation learning, and CodeT5 [24] which proposes an identifier-aware pre-training task to improve embedding distinction. ContraBin aims similar goals with these large-scale language representation models but focuses on a more effective and efficient representation for binary code in terms of better semantics
@@
Contrastive learning As a recent emerging research direction, contrastive learning can enhance the performance of many computer vision tasks by contrasting samples against each other to learn common properties [50, 51]. For example, MoCo [52] proposes a dynamic dictionary to facilitate contrastive learning, SimSiam[53] introduces a Siamese network and stop gradient scheme to prevent collapsing in optimization, and CLIP [19] integrates visual concepts and raw text about images to provide much broader source of supervision in multi-modal contrastive learning. As for program understanding, ContraCode [54] is the first work using contrastive learning in source code analysis by code augmentation and comparison. We design ContraBin based on the contrastive learning framework which simultaneously and gradually learns from binary code, source code, and comments within the same program.

### v5

@@ 4 Experimental Results
Figure 6: A t-SNE visualization of binary code embeddings from the AnghaBench test set. Each plot shows the overall embedding space (blue dots) from a different model. The highlighted red and orange dots represent two specific, semantically similar functions (process_record_keymap and process_record_user) to illustrate semantic alignment. Baseline models like RoBERTa place these similar functions far apart and produce a scattered overall distribution. In contrast, both ContraBin (PCL) and the final ContraBin model cluster the similar functions closely and organize the overall embedding space into more structured regions, demonstrating an improved ability to capture semantic relationships.
@@ 4.1 Embedding Analysis
To ensure a fair and direct comparison, all baseline models presented in our evaluation were subjected to the exact same fine-tuning protocol as ContraBin. This standardized procedure includes the use of the identical triplet dataset, data splits (training, validation, and test), optimizer (AdamW), learning rate schedule, batch sizes, and early-stopping criterion. By maintaining a consistent experimental setup across all models, we can attribute performance differences primarily to the quality of the pre-trained embeddings themselves.
@@
Table 3: Quantitative evaluation of ContraBin on binary functional algorithm classification of POJ104 using mean of average precision (MAP), recall (MAR) and F1 score (MAF1).
@@
RoBERTa (Liu et al., 2019) 38.29 28.57 32.72
CodeBERT (Feng et al., 2020) 39.99 (+4.44%) 29.94 (+4.80%) 34.24 (+4.64%)
GraphCodeBERT (Guo et al., 2020) 39.48 (+3.11%) 29.38 (+2.95%) 33.69 (+2.96%)
@@
Table 4: Quantitative evaluation of ContraBin and multiple pre-trained methods on binary function name recovery of DIRE using accuracy, mean of average precision (MAP), mean of average recall (MAR), and mean of average F1 score (MAF) as four key criteria.
@@
RoBERTa (Liu et al., 2019) 29.41 28.41 25.00 26.59
CodeBERT (Feng et al., 2020) 24.94 (-15.20%) 23.25 (-18.15%) 20.82 (-16.70%) 21.97 (-17.39%)
GraphCodeBERT (Guo et al., 2020) 25.95 (-11.76%) 27.59 (-2.87%) 22.52 (-9.91%) 24.80 (-6.74%)
@@
Our evaluation protocol for all downstream tasks strictly adheres to the standards set by the official CodeXGlue benchmark. This includes the mandated use of its specified default random seed of 123456 for all experiments. This decision is crucial for ensuring that our results are fully reproducible and maintain direct comparability with the established body of work evaluated on this framework.
@@ 4.2 RQ1: Binary Code Analysis
After analyzing embedding of binary code using all pre-trained models, we start qualitative binary code analysis of ContraBin by two downstream tasks: algorithmic functionality classification and binary functiona name recovery.
@@
To evaluate the performance of ContraBin in this downstream task, we adopt mean of average precision (MAP), mean of average recall (MAR), and mean of average F1 score (MAF) as three key criteria. We consider only C code that can be compiled to binaries — while most programs compiled successfully, we exclude those that could not be compiled, and report MAP, MAR, and MAF across all 104 classes. For this task and the three subsequent tasks, we used the results of RoBERTa as a baseline for comparison.
Table 3 presents the quantitative evaluation of ContraBin against several pre-trained methods. CodeBERT and GraphCodeBERT shows comparable performance to RoBERTa on all three metrics, with a performance improvement of less than 5%. However, ContraBin demonstrates a substantial improvement in MAP, MAR, and MAF1 by 14.34%, 18.55%, and 16.71%, respectively. We note that the performance of ContraBin (PCL) is lower than that of RoBERTa. This can be partially attributed to the use of contrastive learning over three different modalities, which can result in the binary code embedding becoming overfitted to the exact embedding of either source code or comments, making the model less effective.
@@
The overall performance, as shown in Table 4, revealed that fine-tuning CodeBERT and GraphCodeBERT resulted in a decrease in performance, making them less effective than the baseline RoBERTa model in all metrics by up to approximately 20%. This demonstrated that current models are unable to generalize when the domain of downstream applications for binary code analysis shifts and may even become ineffective. On the other hand, ContraBin still demonstrated considerable improvement over all three metrics by 13.6%, 8.39%, and 10.58%, respectively, demonstrating the generalizability of our pretrained model on different downstream tasks.
Table 5: Quantitative evaluation of ContraBin and multiple pre-trained methods on binary code summarization of AnghaBench using BLEU-4, GLEU-4, ROUGE-2, and exact match (xMatch). Performance differences are calculated relative to GraphCodeBERT.
Approaches BLEU-4 GLEU-4 ROUGE-2 xMatch
RoBERTa (Liu et al., 2019) 24.30 (-25.05%) 25.55 (-22.45%) 30.16 (-23.52%) 4.95 (-37.41%)
CodeBERT (Feng et al., 2020) 32.07 (-1.11%) 32.80 (-0.43%) 38.83 (-1.57%) 7.91 (0.00%)
GraphCodeBERT (Guo et al., 2020) 32.43 (Base) 32.94 (Base) 39.45 (Base) 7.91 (Base)
@@
Table 6: Quantitative evaluation of ContraBin on binary reverse engineering of AnghaBench using BLEU-4, GLEU-4 and exact match (xMatch).
Approaches BLEU-4 GLEU-4 xMatch
RoBERTa (Liu et al., 2019) 69.72 68.29 23.52
CodeBERT (Feng et al., 2020) 70.37 (+0.94%) 67.27 (-1.50%) 23.41 (-0.47%)
GraphCodeBERT (Guo et al., 2020) 71.56 (+2.64%) 69.88 (+2.32%) 24.73 (+5.14%)
@@ 4.3 RQ2: Binary Code Comprehension
Result for AnghaBench (Summarization) For the binary code summarization task, we evaluate all models using four widely adopted metrics: BLEU-4 (Papineni et al., 2002), GLEU-4 (Mutton et al., 2007), ROUGE-2 (Lin, 2004), and Exact Match (xMatch). As shown in Table 5, ContraBin achieves consistent improvements across all metrics compared to the baseline method, GraphCodeBERT. Specifically, ContraBin shows improvements of 5.95% in BLEU-4, 5.71% in GLEU-4, 4.43% in ROUGE-2, and 18.08% in xMatch over GraphCodeBERT, showcasing its superior capability in binary code summarization.
To further highlight the effectiveness of ContraBin, we present a case study in Figure 7. In the first example, ContraBin successfully captures the functionality of the binary program, while other models diverge to irrelevant objectives, such as "cpu," "file," or "lock." In the second example, other models struggle with summarizing longer binary code. For instance, ContraBin (PCL) generates incomplete sentences, whereas RoBERTa, CodeBERT, and GraphCodeBERT produce either incorrect or overly general summaries. These cases demonstrate ContraBin’s ability to maintain long-term semantic consistency and robustness in binary code summarization, enabled by its novel binary embeddings.
Result for AnghaBench (Reverse Engineering) For the binary code reverse engineering task, we use similar metrics, except ROUGE, which is designed for summarization only, to evaluate the performance of ContraBin on direct code translation between source code and binary code. As shown in Table 6, ContraBin improves the baseline method by 2.04% in BLEU, 3.10% in GLEU, and 7.01% in xMatch. Compared with GraphCodeBERT, which outperforms ContraBin in BLEU, ContraBin outperforms on the other two metrics (GLEU and xMatch).
@@ 5 Ablation Studies and Model Insights
Table 7: Ablation studies of ContraBin on binary functional algorithm classification of POJ104. We report mean average precision (MAP), mean average recall (MAR), and mean average F1 score (MAF1). Absolute scores are shown, with differences from the full ContraBin baseline reported in parentheses (pt). Note: Human-written comments provide minimal discriminative signal, resulting in very low absolute performance.
Ablation Variant MAP MAR MAF1
Human-written comments 3.11 (-40.67) 3.01 (-30.86) 3.06 (-35.13)
Comment removal 42.76 (-1.02) 33.23 (-0.64) 37.40 (-0.79)
Anchor removal 40.96 (-2.82) 31.07 (-2.80) 35.34 (-2.85)
Interpolation removal 37.03 (-6.75) 26.99 (-6.88) 31.22 (-6.97)
ContraBin (Full model) 43.78 33.87 38.19
@@ 5.1 Ablation Studies
We assess each component of our model through ablation studies. Specifically, we remove three conditions to test the efficiency of each design element: (1) We remove comments as one of the modalities during training, (2) we replace the anchored encoder with a learnable one so that it can be updated during training, and (3) we eliminate both linear and non-linear interpolation from our model design. For the pretraining process, we adhere to the original hyperparameters, training the first and second ablations for 30 epochs and the third ablation for 10 epochs. For the third ablation, we ensure that the model is converged by the end of pretraining. We test the pretrained models on the binary functional algorithm classification of POJ-104, using the same hyperparameters. The results are shown in Table 7.
@@ 5.2 Comparison of Human-Written and LLM-Generated Comments in Pretraining
Performance comparison The experiments reveal a striking decrease in performance when ContraBin is trained with human-written comments compared to LLM-generated comments. Specifically, Table 7 shows that the use of human-written comments results in MAP dropping from 43.78 to 3.11, MAR from 33.87 to 3.01, and MAF1 from 38.19 to 3.06. These reductions of 40.67 pt, 30.86 pt, and 35.13 pt, respectively, highlight the stark contrast in effectiveness between the two types of comments.
@@ 5.5 Extensibility to Contemporary LLMs
The evolving landscape of LLMs necessitates the adaptation of existing methodologies to newer, more advanced architectures. As LLMs like T5 (Raffel et al., 2020), CodeT5 (Wang et al., 2021) and its derivatives become increasingly prominent, it is crucial to ensure that techniques initially developed for models like RoBERTa can be effectively extended to these contemporary frameworks. This section outlines the key adaptations required to transition from RoBERTa’s encoder-only architecture to T5’s encoder-decoder structure.
@@
Table 8: Performance Comparison of ContraBin and CodeT5 with T5 on Binary Functional Algorithm Classification for POJ104 (MAP, MAR, MAF1).
@@
As shown in Table 8, the ContraBin (CodeT5) model with our pretraining method achieves significant improvements over CodeT5. Specifically, our model achieves a MAP of 30.1, which is a 4.45% improvement over CodeT5, a MAR of 22.16 (+2.07%), and a MAF1 of 25.53 (+3.11%). These results underscore the effectiveness of our training methodology in enhancing the model’s ability to comprehend and process binary code, demonstrating the adaptability and strength of the T5 architecture.
@@ 6 Related Work
Binary code analysis Binary code analysis plays an important role in the bigger domain of software analysis and maintenance. For example, tracing execution can help the analysis of the functionality algorithms from binaries (Pierce & Mudge, 1994), reusing profile information can speed up the similarity comparison of frequently executed core code (Wang et al., 2000), incorporating sequences of system calls can help detect software vulnerabilities (Giffin et al., 2004), and combining machine and binary-interface descriptions can assist software reverse engineering (Cifuentes et al., 1999). Traditionally, in programming language and software engineering, researchers have presented various work based on static and dynamic program analysis techniques. For instance, BYTEWEIGHT (Bao et al., 2014) automatically classifies algorithms by functionality from binaries, BINSEC (Djoudi & Bardin, 2015; David et al., 2016) formalizes low-level regions of code to extract semantics on the programs, and BinSide (Aslanyan et al., 2020) uses intermediate representations to conduct cross-platform static analysis.
Neural code models In recent years, neural code models have attracted the attention of researchers in software engineering and security. With the assistance of AI (Vaswani et al., 2017), it can improve many source code analysis tasks, such as neural clone detection, neural code similarity comparison, neural code search, etc (Ben-Nun et al., 2018; Shi et al., 2019; Zeng et al., 2021; Ye et al., 2020; Zhang et al., 2022).
@@
For example, SySeVR combines syntax, semantics, and neural vector representation to detect software vulnerabilities (Li et al., 2021), InnerEye adopts techniques in Natural Language Processing (NLP) to compare between samples of binary code (Zuo et al., 2019), GNN-BPE (Guo et al., 2022b) applies Graph Neural Networks (GNN) on binary code by fusing the semantics of Control Flow Graphs (CFG), Data Flow Graphs (DFG), and call graphs, and Bin2Vec (Arakelyan et al., 2021) learns binary code representation via Graph Convolutional Networks (GCN). Compiler chain detection has also emerged as a promising application of machine learning in binary analysis, enabling the identification of the toolchains used to generate binary executables and aiding tasks like provenance analysis and malware detection (Chen et al., 2022a; De Blaere et al., 2023). Additionally, the survey by Marcelli et al. (Marcelli et al., 2022) provides a detailed review of how machine learning techniques are solving the binary function similarity problem, which complements the above-discussed approaches. However, these methods involve complex aggregation of additional binary code representations and have limited generalizability across tasks.
@@
Large-scale pre-trained embeddings With the recent success of large-scale pre-trained embedding in AI, research has been conducted on applying similar approaches to code analysis (Ben-Nun et al., 2018; Shi et al., 2022; Jiang et al., 2022; Bertolotti & Cazzola, 2022). In particular, RoBERTa (Liu et al., 2019) serves as the baseline, CodeBERT (Feng et al., 2020) incorporate code generation and identification into training process, GraphCodeBERT (Guo et al., 2020) combines Abstract Syntax Tree (AST) to further improve embedding quality. There are also many recent models designed for source code understanding, including CodeTrans (Elnaggar et al., 2021) and CoTexT (Phan et al., 2021) that incorporate different language modalities and multi sub-tasks for better code representation learning, and CodeT5 (Wang et al., 2021) which proposes an identifier-aware pre-training task to improve embedding distinction. ContraBin aims similar goals with these large-scale language representation models but focuses on a more effective and efficient representation for binary code in terms of better semantics
@@
Contrastive learning As a recent emerging research direction, contrastive learning can enhance the performance of many computer vision tasks by contrasting samples against each other to learn common properties (Tian et al., 2020; Xiao et al., 2020). For example, MoCo (He et al., 2020) proposes a dynamic dictionary to facilitate contrastive learning, SimSiam(Chen & He, 2021) introduces a Siamese network and stop gradient scheme to prevent collapsing in optimization, and CLIP (Radford et al., 2021) integrates visual concepts and raw text about images to provide much broader source of supervision in multi-modal contrastive learning. As for program understanding, ContraCode (Jain & Jain, 2021) is the first work using contrastive learning in source code analysis by code augmentation and comparison. We design ContraBin based on the contrastive learning framework which simultaneously and gradually learns from binary code, source code, and comments within the same program.